# The Peak Frequency and Luminosity of Synchrotron Emitting Shocks: from Non-Relativistic to Ultra-Relativistic Explosions

### v1

@@ 2 The Physical Model
We follow the non-relativistic framework of Margalit & Quataert (2021), extending this model to arbitrary shock velocity. We additionally adopt a line-of-sight (LOS) approximation in which flux is considered to be dominated by material along the LOS to the source. For relativistically expanding sources, emission from different annular regions will reach the observer at different times. A complete calculation of the observed flux in this case therefore necessitates full knowledge of the shock expansion history. This has been performed numerically for the case of GRBs assuming Blandford-McKee expansion (Granot et al., 1999a, b), or as post-processing of hydrodynamic simulations (e.g., van Eerten et al., 2012). Wishing to avoid imposing any specific expansion history (which may differ in different contexts and for different shock velocities), and in the interest of obtaining analytically-tractable results—we adopt the LOS approximation. The Doppler factor is therefore taken to be $\mathcal{D}=\Gamma$, where $\Gamma=1/\sqrt{1-\beta^{2}}$ is the Lorentz factor of post-shock emitting matter with velocity $\beta c$. The lab-frame shock radius can be expressed as
@@
Based on the results of kinetic simulations (e.g., Sironi et al. 2013) we adopt the following fiducial values for the microphysical parameters described above: $\epsilon_{T}=0.4$, $\epsilon_{e}=0.01$ ($\delta=0.025$), $\epsilon_{B}=0.1$, and $p=3$. This implies that 40% of the post-shock energy is carried by thermal electrons, 10% by magnetic fields, and 1% by non-thermal (power-law) accelerated electrons. The remaining $\sim$50of post-shock energy is carried by thermal and non-thermal (cosmic ray) ions. Typical kinetic simulations suggest that $\sim$10of energy goes into cosmic-ray ions and therefore that thermal ions carry $\sim$40of the post-shock energy. This situation implies approximate post-shock equilibration between electrons and ions such that $T_{e}\approx T_{\rm ion}$. Considering lower values of $\epsilon_{T}$ would relax this assumption and generally allow for $T_{e}<T_{\rm ion}$.
@@
From hydrodynamic considerations one expects that $\Delta R^{\prime}\lesssim R/4\Gamma$. We therefore adopt $f=3/16$ as a fiducial value. The LOS averaged emissivity and absorption coefficients, $\left\langle j^{\prime}_{\nu}\right\rangle$, $\left\langle\alpha^{\prime}_{\nu}\right\rangle$, are described in Margalit & Quataert (2021) and include the contribution of both thermal and non-thermal (power-law) electrons. In the non-relativistic limit and with $f=3/4$, Equation (2) reduces to Equation (21) of Margalit & Quataert (2021). In cases where only power-law electrons contribute to observed emission, it is also consistent with the estimates of Chevalier (1998) in the non-relativistic limit, and with Sari et al. (1998) in the ultra-relativistic optically-thin limit. As previously mentioned, in the relativistic case one must generally integrate the specific intensity over a surface from which all emitted photons arrive at the same time (Granot et al., 1999a, b). In Equation (2) we have instead adopted the LOS approximation and neglected this aspect for the sake of analytic tractability. The impact of this assumption will be investigated in future work.
@@
For reasons that will later become apparent, this parameterization is useful because the mass-loss parameter depends on the peak frequency and observed time only through the combination $\nu_{\rm pk}t$. Note that Equation (4) is merely a parameterization choice which replaces $n\leftrightarrow\dot{M}/v_{\rm w}$. In particular, our results are not limited to the restricted case of a steady-state wind medium, and can be applied to arbitrary density profiles.222 One caveat to this statement is that emission is assumed to be dominated by electrons that have been swept-up by the most recent shock-doubling time. This typically requires that most of the swept-up material must be accumulated at $r\sim R$, and for a power-law density gradient $n\propto r^{-k}$, implies that $k<3$. The mass-loss parameter should therefore be viewed as only a proxy for the ambient density at the current shock location.
@@ 3 The Spectral Energy Distribution
The unabsorbed SED is proportional to $j_{\nu}(\nu)$ and is intrinsically peaked at frequency $\nu_{\Theta}=3\Gamma\Theta^{2}eB/4\pi m_{e}c$ (Margalit & Quataert, 2021), the frequency at which a typical thermal electron radiates most of its synchrotron power. The observed SED however will be modified due to SSA and thus does not track the emissivity at frequencies below the SSA frequency, $\nu_{\rm a}$. Two categories therefore emerge when considering the observed peak of the SED. In the first case, the intrinsic peak of the emissivity at $\nu\simeq\nu_{\Theta}$ is hidden due to SSA, and the observed peak of the SED occurs instead at the SSA frequency $\nu_{\rm a}$. This happens when $\nu_{\Theta}<\nu_{\rm a}$, that is—when the SSA optical depth at frequency $\nu_{\Theta}$ is $\tau_{\Theta}>1$. The second case occurs when $\tau_{\Theta}<1$ and the SSA frequency falls below the thermal frequency, $\nu_{\rm a}<\nu_{\Theta}$. In this case the intrinsic peak of the emissivity is unabsorbed and corresponds to the observed peak of the SED ($\nu_{\rm pk}\simeq\nu_{\Theta}$). The SSA optical depth is a strong function of shock velocity (Margalit & Quataert, 2021), and we show in subsequent sections that the two scenarios described above are delineated by whether the velocity is smaller or larger than $\left(\Gamma\beta\right)_{\rm sh}\sim 1$. The SSA SED can further be divided into two subcases depending on whether thermal electrons or power-law electrons dominate emission at frequency $\nu_{\rm pk}$. As discussed above, these two cases correspond to shock velocities above or below $\left(\Gamma\beta\right)_{\rm sh}\sim 0.3$, respectively.
@@ 4.1 Power-law Electrons
We note that the numerical prefactors in Equations (4.1,4.1) differ slightly from the corresponding expressions in Chevalier (1998). Specifically, the shock velocity implied by Equation (4.1) is $\sim$30larger than would be inferred from Equation (13) in Chevalier (1998). This is due to two factors. First, Chevalier (1998) expresses results in terms of a “peak” luminosity/frequency that is defined by the intersection of the optically-thin and optically-thick power-law limits, instead of using the true peak values as implied by Equation (2). This method overestimates the peak luminosity and underestimates peak frequency.333 More precisely, the “peak” frequency $\nu_{*}$ as defined in Chevalier (1998) is related to the true peak frequency $\nu_{\rm pk}$ via $\nu_{*}=\nu_{\rm pk}\tau_{\rm pk}^{2/(p+4)}$. The peak luminosity is similarly given by $L_{\nu_{*}}=L_{\nu_{\rm pk}}\tau_{\rm pk}^{5/(p+4)}/(1-e^{-\tau_{\rm pk}})$. The optical depth at peak $\tau_{\rm pk}$ can be found by numerically solving the equation $e^{-\tau_{\rm pk}}\left[(p+4)\tau_{\rm pk}+5\right]=5$ (e.g., Ghisellini 2013 § A.5). For $p=3$, $\tau_{\rm pk}\simeq 0.639$ and therefore $\nu_{*}\simeq 0.88\nu_{\rm pk}$ and $L_{\nu_{*}}\simeq 1.54L_{\nu_{\rm pk}}$. Second, a small multiplicative difference between the two prescriptions arises due to different treatments of the electron pitch angle. Chevalier (1998) sets $\sin\theta=1$ for simplicity whereas we assume a more realistic scenario where emitting electrons are randomly oriented (see Margalit & Quataert 2021). Combining these effects, one can show that the above expressions are consistent with the Chevalier (1998) results if one replaces $L_{\nu_{\rm pk}}\to 2.02L_{\nu_{\rm pk}}$ and $\nu_{\rm pk}\to 1.01\nu_{\rm pk}$ in the equations of Chevalier (1998).
@@ 5 Optically-Thin Peak: $\left(\Gamma\beta\right)_{\rm sh}\gg$$\gtrsim 1$
These expressions are relevant for ultra-relativistic shocks such as in GRBs. Emission by thermal electrons in GRBs has been investigated by several studies in recent years (Giannios & Spitkovsky, 2009; Warren et al., 2017; Ressler & Laskar, 2017; Warren et al., 2018; Samuelsson et al., 2020; Warren et al., 2022). Our present results add to this growing literature by providing analytic expressions that allow one to infer the shock properties from the (thermal-electron-dominated) SED peak, if a clear peak in the SED is detected. This regime may also be relevant in other astrophysical contexts as well. For example the ‘synchrotron-maser’ model for fast radio bursts (FRBs) invokes magnetized ultra-relativistic shocks as a mechanism for powering FRBs. A prediction of this model is that FRBs will be accompanied by high-energy synchrotron emission from thermal electrons behind the shock (Lyubarsky, 2014; Beloborodov, 2017; Metzger et al., 2019; Beloborodov, 2020; Margalit et al., 2020b, a). Our present analysis would be suitable for describing such emission, albeit only if emitting electrons are slow-cooling (in the FRB context it is more likely that emission is instead fast-cooling; e.g., Metzger et al. 2019). To date, only a single FRB has an observed high-frequency counterpart that may be attributed to this model (the Galactic FRB 20200428; Margalit et al. 2020a). However future Galactic or local-Universe FRBs may increase the number of such detections.
@@ 6.1 Energy
when thermal electrons are responsible for the observed emission at frequency $\nu_{\rm pk}$. In the regime where it is appropriate ($L_{\nu_{\rm pk}}>L_{\rm th}$), Equation (17) gives a lower value of the post-shock energy than would have otherwise been inferred using the standard Chevalier (1998) model (Equation 16). In particular, the ratio of these inferred energies can be expressed as $U[\text{thermal }e^{-}]/U[\text{power-law }e^{-}]\sim\left(L_{\nu_{\rm pk}}/L_{\rm th}\right)^{-23/19}$, with only very weak dependence on other parameters.444 $U[\text{thermal }e^{-}]/U[\text{power-law }e^{-}]\propto\left(\epsilon_{B}/f\nu_{\rm pk}\bar{t}\right)^{0.08}\epsilon_{T}^{0.03}\epsilon_{e}^{-0.07}$ This illustrates how the inferred energy drops when one accounts for thermal electrons in the regime $L_{\nu_{\rm pk}}\gg L_{\rm th}$ where they are important.
@@
Figure 5 illustrates this point. Colored symbols show the inferred energy and shock velocity for different transients as implied by our model. Red stars show the currently known sample of FBOTs, while differently colored open circles show various SNe. Each colored point is connected to a corresponding smaller grey symbol that shows the properties that would have been inferred for the same event if relativistic corrections and thermal electrons would have been neglected (in some cases the points overlap such that the grey symbols are hidden behind the colored ones). Events that fall to the left of this plot (lower velocities) satisfy $L_{\nu_{\rm pk}}<L_{\rm th}$ such that their observed emission is dominated by power-law electrons and their SED will be similar to that shown in Figure 1 panel (a). Conversely, events that fall to the right of the leftmost vertical dashed-black curve have their peak emission dominated by thermal electrons ($L_{\nu_{\rm pk}}>L_{\rm th}$).555 The exact velocity at which $L_{\nu_{\rm pk}}=L_{\rm th}$ depends on $\nu_{\rm pk}t$ (albeit weakly; Equation 8) and will therefore vary for different events, see Figure 2. In Figure 5 we plot this transition for a representative $\nu_{\rm pk}\bar{t}=5\,{\rm GHz}\times 100\,{\rm day}$. Figure 5 shows that while the non-relativistic power-law-electron framework of Chevalier (1998) works well for events where the shock velocity is $\lesssim 0.2c$, it is inaccurate for trans-relativistic events with higher velocity. In particular, the shock velocity and energy are overestimated by this model (compare grey points with their associated colored symbols), the latter by as much as two orders of magnitude. This discrepancy becomes more acute as one moves towards higher shock velocities and further away from the $L_{\nu_{\rm pk}}=L_{\rm th}$ line. This highlights the importance of considering relativistic corrections and thermal electrons when modeling such events.
We note that the discussion above applies to the physically-motivated scenario where $\epsilon_{e}\ll\epsilon_{T}\lesssim 1$ (that is, $\delta\ll 1$). A minimum on the explosion energy is often estimated by adopting ‘equipartition’ microphysical parameters instead, $\epsilon_{e}\sim\epsilon_{B}\sim\mathcal{O}(1)$. One can immediately see from Equation (16) that such values of $\epsilon_{e}$ and $\epsilon_{B}$ indeed minimize the inferred energy. In this scenario $\epsilon_{e}\sim\epsilon_{T}$ (i.e., $\delta\sim 1$), so that $L_{\rm th}$ is very large and there is little to no phase space in which thermal electrons dominate peak emission. That is, if power-law electrons are assumed to be in equipartition with magnetic fields and thermal electrons + ions, then $L_{\rm th}\sim L_{\rm crit}$ such that $L_{\nu_{\rm pk}}$ is almost always smaller than $L_{\rm th}$, and therefore Equation (16)—and not (17)—is appropriate for estimating the energy. In other words, in the equipartition scenario thermal electrons are almost never relevant, even if they are present. This implies that the standard minimum-energy estimates, which neglect thermal electrons entirely, are still valid lower-limits on $U$. The reduction in inferred energies when including thermal electrons that is discussed in previous paragraphs applies to the physically motivated case where $\epsilon_{e}\ll 1$ rather than equipartition. This would alleviate energetic constraints for more reasonable expected values of $\epsilon_{e}$ and $\epsilon_{B}$. We also note that an equipartition argument holds true also in the case where thermal electrons are included and dominant over power-law electrons ($\epsilon_{e}\ll\epsilon_{T}$, $L_{\nu_{\rm pk}}>L_{\rm th}$). This can be derived from Equation (17), which again shows that $U$ is minimized when $\epsilon_{T}\sim\epsilon_{B}\sim\mathcal{O}(1)$.
Although it is reasonable to expect that $\epsilon_{e}\ll\epsilon_{T}$, the efficiency of electron-ion equilibration is not fully understood and therefore the extent to which $\epsilon_{T}\sim 1$ is somewhat uncertain. Particle-in-cell simulations tend to support the idea that plasma processes can mediate efficient energy transfer between ions and electrons (Coulomb collisions are typically ineffective at this), however this may not necessarily lead to full equilibration (where the temperature of ions and electrons are the same), and the details are still debated (e.g., Sironi et al., 2013; Tran & Sironi, 2020; Raymond et al., 2023; Groselj et al., 2024). [^6] Note that a strict lower limit $\epsilon_{T}>\mu_{e}m_{e}/\mu m_{p}\simeq 10^{-3}$ arises from considering the case where electrons and ions do not exchange energy at all.

### v2

@@ 2 The Physical Model
We follow the non-relativistic framework of Margalit & Quataert (2021), extending this model to arbitrary shock velocity. We additionally adopt an effective line-of-sight (LOS) approximation in which flux is considered to be dominated by material at some representative point within the source. For relativistically expanding sources, emission from different regions will reach the observer at different times. A complete calculation of the observed flux in this case therefore necessitates a numerical approach and requires knowledge of the shock expansion history. This has been performed numerically for the case of GRBs assuming Blandford-McKee expansion (Granot et al., 1999a, b), or as post-processing of hydrodynamic simulations (e.g., van Eerten et al., 2012). Wishing to avoid imposing any specific expansion history (which may differ in different contexts and for different shock velocities), and in the interest of obtaining analytically-tractable results—we adopt the LOS approximation. As part of this approximation we take the effective Doppler factor to be $\mathcal{D}=\Gamma$, where $\Gamma=1/\sqrt{1-\beta^{2}}$ is the Lorentz factor of post-shock emitting matter with velocity $\beta c$. We also assume that the shock radius (as measured in the lab-frame) is
@@
Based on the results of kinetic simulations (e.g., Sironi et al. 2013) we adopt the following fiducial values for the microphysical parameters described above: $\epsilon_{T}=0.4$, $\epsilon_{e}=0.01$ ($\delta=0.025$), $\epsilon_{B}=0.1$, and $p=3$. This implies that 40% of the post-shock energy is carried by thermal electrons, 10% by magnetic fields, and 1% by non-thermal (power-law) accelerated electrons. The remaining $\sim$50of post-shock energy is carried by thermal and non-thermal (cosmic ray) ions. Typical kinetic simulations suggest that $\sim$10of energy goes into cosmic-ray ions and therefore that thermal ions carry $\sim$40of the post-shock energy. This situation implies approximate post-shock equilibration between electrons and ions such that $T_{e}\sim T_{\rm ion}$. Considering lower values of $\epsilon_{T}$ would relax this assumption and generally allow for $T_{e}<T_{\rm ion}$. We discuss these assumptions further in §6.2.
@@
From hydrodynamic considerations one expects that $\Delta R^{\prime}\lesssim R/4\Gamma$. Furthermore, requiring that the number of electrons in the circumstellar medium that have been swept up by the shock equals the number of electrons within the post-shock shell (assuming a constant density shell of emitting matter) leads to $f\leq{3}/{\left[16(3-k)\right]}$ where $30\leq k<$ 3 is the slope of the circumstellar density profile, such that $n(r)\propto r^{-k}$. A larger value of $f$ would imply that the number of particles within the post-shock shell is larger than the number of swept-up particles, which is not possible. Smaller values of $f$ are permissible if one assumes that only a fraction of the total swept-up material contributes to synchrotron emission (and absorption). For our estimates below we adopt $f=3/16$ as a fiducial value.
In Equation (2) above, the LOS averaged emissivity and absorption coefficients, $\left\langle j^{\prime}_{\nu}\right\rangle$, $\left\langle\alpha^{\prime}_{\nu}\right\rangle$, are given by Margalit & Quataert (2021) and include the contribution of both thermal and non-thermal (power-law) electrons. In the non-relativistic limit and with $f=3/4$, Equation (2) reduces to Equation (21) of Margalit & Quataert (2021). In cases where only power-law electrons contribute to observed emission, it is also consistent with the estimates of Chevalier (1998) in the non-relativistic limit, and with Sari et al. (1998) in the ultra-relativistic optically-thin limit.
As previously mentioned, in the relativistic case one must generally integrate the specific intensity over a region from which all emitted photons arrive at the same time (Granot et al., 1999a, b). In Equation (2) we have instead adopted an effective LOS approximation and neglected this aspect for the sake of analytic tractability. This is similar to the approach taken by Sari et al. (1998). The impact of this assumption will be investigated in future work.222 On a technical note—we have called the formalism above an ‘effective’ LOS approach because, although it does rely on estimating the total flux based on the intensity along some representative ray (that is, emission from some representative region), this region is not precisely along the line-of-sight between the origin of the explosion the observer (otherwise the Doppler factor would be larger by a factor $1+\beta$). This choice is equally valid given that, in any case, the approach is merely an approximation to the more complicated relativistic radiative transfer problem that must be solved. We chose this particular formalism because it coincides with the assumptions of Sari et al. (1998) for GRBs.
@@
For reasons that will later become apparent, this parameterization is useful because the mass-loss parameter depends on the peak frequency and observed time only through the combination $\nu_{\rm pk}t$. Note that Equation (4) is merely a parameterization choice which replaces $n\leftrightarrow\dot{M}/v_{\rm w}$. In particular, our results are not limited to the restricted case of a steady-state wind medium, and can be applied to arbitrary density profiles.333 One caveat to this statement is that emission is assumed to be dominated by electrons that have been swept-up by the most recent shock-doubling time. This typically requires that most of the swept-up material must be accumulated at $r\sim R$, and for a power-law density gradient $n\propto r^{-k}$, implies that $k<3$. The mass-loss parameter should therefore be viewed as only a proxy for the ambient density at the current shock location.
@@ 3 The Spectral Energy Distribution
The unabsorbed SED is proportional to $j_{\nu}(\nu)$ and is intrinsically peaked at frequency $\nu_{\Theta}^{\prime}=3\Theta^{2}eB/4\pi m_{e}c$ (Margalit & Quataert, 2021) in the rest-frame of emitting matter, or $\nu_{\Theta}=\mathcal{D}\nu_{\Theta}^{\prime}$ in the observer frame. This corresponds to, the frequency at which a typical thermal electron radiates most of its synchrotron power. The observed SED however will be modified due to SSA and thus does not track the emissivity at frequencies below the SSA frequency, $\nu_{\rm a}$. Two categories therefore emerge when considering the observed peak of the SED. In the first case, the intrinsic peak of the emissivity at $\nu\simeq\nu_{\Theta}$ is hidden due to SSA, and the observed peak of the SED occurs instead at the SSA frequency $\nu_{\rm a}$. This happens when $\nu_{\Theta}<\nu_{\rm a}$, that is—when the SSA optical depth at frequency $\nu_{\Theta}$ is $\tau_{\Theta}>1$. The second case occurs when $\tau_{\Theta}<1$ and the SSA frequency falls below the thermal frequency, $\nu_{\rm a}<\nu_{\Theta}$. In this case the intrinsic peak of the emissivity is unabsorbed and corresponds to the observed peak of the SED ($\nu_{\rm pk}\simeq\nu_{\Theta}$). The SSA optical depth is a strong function of shock velocity (Margalit & Quataert, 2021), and we show in subsequent sections that the two scenarios described above are delineated by whether the velocity is smaller or larger than $\left(\Gamma\beta\right)_{\rm sh}\sim 1$. The SSA SED can further be divided into two subcases depending on whether thermal electrons or power-law electrons dominate emission at frequency $\nu_{\rm pk}$. As discussed above, these two cases correspond to shock velocities above or below $\left(\Gamma\beta\right)_{\rm sh}\sim 0.3$, respectively.
@@ 4.1 Power-law Electrons
We note that the numerical prefactors in Equations (4.1,4.1) differ slightly from the corresponding expressions in Chevalier (1998). Specifically, the shock velocity implied by Equation (4.1) is $\sim$30larger than would be inferred from Equation (13) in Chevalier (1998). This is due to two factors. First, Chevalier (1998) expresses results in terms of a “peak” luminosity/frequency that is defined by the intersection of the optically-thin and optically-thick power-law limits, instead of using the true peak values as implied by Equation (2). This method overestimates the peak luminosity and underestimates peak frequency.444 More precisely, the “peak” frequency $\nu_{*}$ as defined in Chevalier (1998) is related to the true peak frequency $\nu_{\rm pk}$ via $\nu_{*}=\nu_{\rm pk}\tau_{\rm pk}^{2/(p+4)}$. The peak luminosity is similarly given by $L_{\nu_{*}}=L_{\nu_{\rm pk}}\tau_{\rm pk}^{5/(p+4)}/(1-e^{-\tau_{\rm pk}})$. The optical depth at peak $\tau_{\rm pk}$ can be found by numerically solving the equation $e^{-\tau_{\rm pk}}\left[(p+4)\tau_{\rm pk}+5\right]=5$ (e.g., Ghisellini 2013 § A.5). For $p=3$, $\tau_{\rm pk}\simeq 0.639$ and therefore $\nu_{*}\simeq 0.88\nu_{\rm pk}$ and $L_{\nu_{*}}\simeq 1.54L_{\nu_{\rm pk}}$. Second, a small multiplicative difference between the two prescriptions arises due to different treatments of the electron pitch angle. Chevalier (1998) sets $\sin\theta=1$ for simplicity whereas we assume a more realistic scenario where emitting electrons are randomly oriented (see Margalit & Quataert 2021). Combining these effects, one can show that the above expressions are consistent with the Chevalier (1998) results if one replaces $L_{\nu_{\rm pk}}\to 2.02L_{\nu_{\rm pk}}$ and $\nu_{\rm pk}\to 1.01\nu_{\rm pk}$ in the equations of Chevalier (1998).
@@ 5 Optically-Thin Peak: $\left(\Gamma\beta\right)_{\rm sh}\gg$$\gtrsim 1$
These expressions are relevant for ultra-relativistic shocks such as in GRBs. Emission by thermal electrons in GRBs has been investigated by several studies in recent years (Giannios & Spitkovsky, 2009; Warren et al., 2017; Ressler & Laskar, 2017; Warren et al., 2018; Samuelsson et al., 2020; Warren et al., 2022). Our present results add to this growing literature by providing analytic expressions that allow one to infer the shock properties from the (thermal-electron-dominated) SED peak, if a clear peak in the SED is detected. In practice however there are very few GRBs where a clear spectral peak in the broadband SED is observed. Additionally, the expressions derived here are only correct before the jet-break time (because of our assumption of spherical symmetry) and in the slow-cooling regime. These aspects limit the range of applicability of these results to GRB afterglows. An extension of this model to include these effects would be straightforward, but is outside the scope of our present work.
The ultra-relativistic results (Equations 13,5) may also be relevant in other astrophysical contexts apart from GRBs. For example the ‘synchrotron-maser’ model for fast radio bursts (FRBs) invokes magnetized ultra-relativistic shocks as a mechanism for powering FRBs. A prediction of this model is that FRBs will be accompanied by high-energy synchrotron emission from thermal electrons behind the shock (Lyubarsky, 2014; Beloborodov, 2017; Metzger et al., 2019; Beloborodov, 2020; Margalit et al., 2020b, a). Our present analysis would be suitable for describing such emission, albeit only if emitting electrons are slow-cooling (in the FRB context it is more likely that emission is instead fast-cooling; e.g., Metzger et al. 2019). To date, only a single FRB has an observed high-frequency counterpart that may be attributed to this model (the Galactic FRB 20200428; Margalit et al. 2020a). However future Galactic or local-Universe FRBs may increase this sample.
@@ 6.1 Energy
when thermal electrons are responsible for the observed emission at frequency $\nu_{\rm pk}$. In the regime where it is appropriate ($L_{\nu_{\rm pk}}>L_{\rm th}$), Equation (17) gives a lower value of the post-shock energy than would have otherwise been inferred using the standard Chevalier (1998) model (Equation 16). In particular, the ratio of these inferred energies can be expressed as $U[\text{thermal }e^{-}]/U[\text{power-law }e^{-}]\sim\left(L_{\nu_{\rm pk}}/L_{\rm th}\right)^{-23/19}$, with only very weak dependence on other parameters.555 $U[\text{thermal }e^{-}]/U[\text{power-law }e^{-}]\propto\left(\epsilon_{B}/f\nu_{\rm pk}\bar{t}\right)^{0.08}\epsilon_{T}^{0.03}\epsilon_{e}^{-0.07}$ This illustrates how the inferred energy drops when one accounts for thermal electrons in the regime $L_{\nu_{\rm pk}}\gg L_{\rm th}$ where they are important.
@@
Figure 5 illustrates this point. Colored symbols show the inferred energy and shock velocity for different transients as implied by our model. Red stars show the currently known sample of FBOTs, while differently colored open circles show various SNe. Each colored point is connected to a corresponding smaller grey symbol that shows the properties that would have been inferred for the same event if relativistic corrections and thermal electrons would have been neglected (in some cases the points overlap such that the grey symbols are hidden behind the colored ones). Events that fall to the left of this plot (lower velocities) satisfy $L_{\nu_{\rm pk}}<L_{\rm th}$ such that their observed emission is dominated by power-law electrons and their SED will be similar to that shown in Figure 1 panel (a). Conversely, events that fall to the right of the leftmost vertical dashed-black curve have their peak emission dominated by thermal electrons ($L_{\nu_{\rm pk}}>L_{\rm th}$).666 The exact velocity at which $L_{\nu_{\rm pk}}=L_{\rm th}$ depends on $\nu_{\rm pk}t$ (albeit weakly; Equation 8) and will therefore vary for different events, see Figure 2. In Figure 5 we plot this transition for a representative $\nu_{\rm pk}\bar{t}=5\,{\rm GHz}\times 100\,{\rm day}$. Figure 5 shows that while the non-relativistic power-law-electron framework of Chevalier (1998) works well for events where the shock velocity is $\lesssim 0.2c$, it is inaccurate for trans-relativistic events with higher velocity. In particular, the shock velocity and energy are overestimated by this model (compare grey points with their associated colored symbols), the latter by as much as two orders of magnitude. This discrepancy becomes more acute as one moves towards higher shock velocities and further away from the $L_{\nu_{\rm pk}}=L_{\rm th}$ line. This highlights the importance of considering relativistic corrections and thermal electrons when modeling such events.